Find the value of K so that:
X^2+Kxy+ 2y^2+3x+5y+2=O 
May represent a line pair.

Grade:10

2 Answers

Arun
3 years ago
Dear student
 
For a pair of straight line
 
Equate abc +2fgh - af² - bg² - ch² = 0
 
Hopw you understand.
 
Thanks and
Aditya Gupta
3 years ago
the condition for a general quad eqn of the form ax^2+by^2+2gx+2fy+2hxy+c=0 to represent a pair of st lines (POSL) is that the determinant should be 0.
here a=1, b= 2, g= 3/2, f= 5/2, h= k/2, c=2, so we form the determinant as follows:
| 1     k/2  3/2|
| k/2  2     5/2|  =  0
|3/2   5/2    2 |
putting this det 0, we get K= 3, 9/2
KINDLY APPROVE :))